# The Arens–Michael envelope of a solvable Lie algebra is a homological epimorphism

### v1

@@ 4. Homological epimorphisms and relatively quasi-free algebras
Let $s\in[0,+\infty]$, $R$ be an associative algebra of countable dimension, $S$ be a $\mathop{\widehat{\otimes}}$-algebra, and $\alpha\!:R\to S$ be a homological epimorphism with dense range. Suppose that $S$ is an $\mathfrak{A}_{s}$$\mathbin{\widehat{\otimes}}$-module algebra, respectively and $R$ is endowed with a derivation is a such way that $\alpha$ becomes a $\mathbb{C}[z]$-module morphism. Then
$\alpha\mathop{\widehat{\#}}\beta\!:R\mathop{\#}\mathbb{C}[z]\to S\mathop{\widehat{\#}}\mathfrak{A}_{s}$
is also a homological epimorphism.
@@
Recall that an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ is said to be relatively quasi-free over $R$ if any admissible singular $R$-extension of $A$ is split; see [28, Definition 7.1] and the discussion therein.
@@
For the proof we need a relative version of Property (UDE). If $A$ is an $R$$\mathop{\widehat{\otimes}}$-algebra and $X$ is a $B$$\mathop{\widehat{\otimes}}$-bimodule, then a derivation $D\!:A\to X$ is called an $R$-derivation if it is an $R$-bimodule morphism (or, equivalently, of left or right $R$-module morphism). The subspace of $\operatorname{Der}(A,X)$ consisting of $R$-derivations is denoted by $\operatorname{Der}_{R}(A,X)$. An $R$-$S$-homomorphism $(f,g)\!:(A,R)\to(B,S)$ induces the linear map
$\operatorname{Der}_{S}(B,X)\to\operatorname{Der}_{R}(A,X)\!:D\mapsto Df;$
see details in [28].
@@
Since each of $f$ and $g$ has dense range, it suffices to verify (4.3) in the case when $s=g(r)$ and $b=f(a)$ for some $r\in R$ and $a\in A$. Then $\eta_{B}(s)b=\eta_{B}(g(r))f(a)=f(\eta_{A}(r))f(a)=f(\eta_{A}(r)a)$ and hence $D^{\prime}(\eta_{B}(s)b)=D(\eta_{A}(r)a)$. Since $D$ is an $R$-derivation, we have $D(\eta_{A}(r)a)=\eta_{A}(r)\cdot D(a)$. On the other hand,
$\eta_{A}(r)\cdot D(a)=f(\eta_{A}(r))\cdot D(a)=\eta_{B}(g(r))\cdot D^{\prime}f(a)=\eta_{B}(s)\cdot D^{\prime}(b).$
@@
The conditions of Proposition 4.8 can be weakened by assuming that $g$ is an epimorphism of $\mathop{\widehat{\otimes}}$-algebras; cf. [28, Proposition 3.8]. But we do nit need this variant.
If $A$ is an $R$$\mathop{\widehat{\otimes}}$-algebra, then a representing object of the functor $\operatorname{Der}_{R}(A,-)$ from $A\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}A$ to the category of sets exists and is denoted by $\Omega^{1}_{R}A$ (it is called the bimodule of relative differential $1$-forms of $A$); for details see [28, p. 82].
@@
where $\mathbf{h}{{}_{B}}(-,-)$ denotes the vector space of $B$$\mathop{\widehat{\otimes}}$-bimodule morphisms (and similarly for $A$); see [28, eq. (7.3) and p. 83]. Also, consider the natural map
@@
Substituting $\Omega_{S}^{1}B$ for $X$, we obtain the morphism
$B\mathop{\widehat{\otimes}}_{A}\Omega^{1}_{R}A\mathop{\widehat{\otimes}}_{A}B\to\Omega_{S}^{1}B$
@@
(cf. [28, Proposition 7.4]) Under the hypotheses of Proposition 4.8, the morphism $B\mathop{\widehat{\otimes}}_{A}\Omega^{1}_{R}A\mathop{\widehat{\otimes}}_{A}B\to\Omega_{S}^{1}B$ defined above is a $B$$\mathop{\widehat{\otimes}}$-bimodule isomorphism.
@@
Let $X$ be an arbitrary $B$$\mathop{\widehat{\otimes}}$-bimodule. Note that the map in (4.5) is always bijective. Also, under the hypotheses of Proposition 4.8, the top arrow in (4.4) is a bijection and so is the bottom arrow.
Thus we have that the map in (4.6) is bijective for every $X$, which implies that $B\mathop{\widehat{\otimes}}_{A}\Omega^{1}_{R}A\mathop{\widehat{\otimes}}_{A}B\to\Omega_{S}^{1}B$ is an isomorphism. ∎
@@
By [28, Proposition 7.3], $A$ is relatively quasi-free over $R$ if and only if $\Omega^{1}_{R}A$ is projective in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(A,R)$. Hence the sequence
@@
is a projective resolution in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(A,R)$. (Here $m$ is the multiplication on $A$ and $j$ is the bimodule morphism corresponding to the inner derivation $-\mathop{\mathrm{ad}}\nolimits_{1\otimes 1}\!:A\to A\mathop{\widehat{\otimes}}_{R}A\!:a\mapsto[1\otimes 1,a]$.)
@@
Since $f$ satisfies Property (UDE), it is an epimorphism. Then the canonical map $B\mathop{\widehat{\otimes}}_{A}B\to B$ is a topological isomorphism. Also, since $g$ is an epimorphism, the canonical map $B\mathop{\widehat{\otimes}}_{R}B\to B\mathop{\widehat{\otimes}}_{S}B$ is a topological isomorphism; for both properties of epimorphisms see [28, Proposition 6.1]. Applying the functor $B\mathop{\widehat{\otimes}}_{A}(-)\mathop{\widehat{\otimes}}_{A}B$ to the sequence in (4.7) and using Proposition 4.10, we get the sequence
$0\leftarrow B\leftarrow B\mathop{\widehat{\otimes}}_{S}B\leftarrow\Omega^{1}_{S}B\leftarrow 0,$
@@
which splits in $B\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}S$ by [28, Proposition 7.2]. Hence it also splits in $S\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}S$. This implies that $f$ is a two-sided relative homological epimorphism; see Definition 4.1. ∎
@@
A projective $(A,R)$$\mathop{\widehat{\otimes}}$-bimodule $P$ satisfies the finiteness condition $(f_{2})$ if it is a retract of a bimodule $A\mathop{\widehat{\otimes}}_{R}M\mathop{\widehat{\otimes}}_{R}A$, where $M\in R\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}R$ is isomorphic to the $R$$\mathop{\widehat{\otimes}}$-bimodule $R^{n}$ for some $n$ [28, Definition 8.3]. (We do not need the alternative condition $(f_{1})$ from [28].) Also, an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ is called an algebra of $(f_{2})$-finite type if in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(A,R)$ it has a finite projective resolution all of whose bimodules satisfy the finiteness condition $(f_{2})$ [28, Definition 8.4].
@@
Recall that we denote by $\beta$ the Hopf $\mathop{\widehat{\otimes}}$-algebra homomorphism $\mathbb{C}[z]\to\mathfrak{A}_{s}$. Put $A=R\mathop{\#}\mathbb{C}[z]$, $B=S\mathop{\widehat{\#}}\mathfrak{A}_{s}$, $g=\alpha$ and $f=\alpha\mathop{\widehat{\#}}\beta$ (the latter is well defined by Lemma 2.3 since $\alpha$ is a $\mathbb{C}[z]$-module morphism). It suffices to check the conditions in Theorem 1.1, i.e.,
@@
Condition (2) follows from Theorem 4.7. Indeed, $\beta$ is a homological epimorphisms by Proposition 4.5. Hence $\alpha$ and $\beta$ satisfy Property (UDE) (the first by the hypothesis and the second by Lemma 4.3). Since the ranges are dense, Theorem 3.4 implies that $f$ also satisfies Property (UDE). Since $R$ is a $\mathbb{C}[z]$-module algebra, we have the action by a derivation. Therefore $R\mathop{\#}\mathbb{C}[z]$ is an Ore extension (see Remark 2.2) and hence it is relatively quasi-free by [28, Proposition 7.9]. Thus all the conditions in Theorem 4.7 hold and therefore $f$ is a two-sided relative homological epimorphism. Finally, note that every two-sided relative homological epimorphism is a one-sided relative homological epimorphism [28, Proposition 6.5].
To verify Condition (3) note that $S\mathop{\widehat{\#}}\mathfrak{A}_{s}$ is isomorphic to $S\mathop{\widehat{\otimes}}\mathfrak{A}_{s}$ as an $S$$\mathop{\widehat{\otimes}}$-module and so it is projective being a free module (see, e.g., [18, Chapter III, Proposition 1.25]). Similarly, $R\mathop{\#}\mathbb{C}[z]$ is a projective $R$-module.
Since $R$ has countable dimension, each Ore extension of $R$ (without twisting) is a relatively $(f_{2})$-quasi-free $R$-algebra [28, Proposition 7.9 and Example 8.1], which in particular implies that it is of $(f_{2})$-finite type [28, Definition 8.5 and Proposition 7.3.]. Thus Condition (4) holds. ∎
@@
For the necessity see [26, Theorem 3.6].
To show the sufficiency consider the decomposition in (2.8). We claim that the truncated homomorphism
@@
Proceeding by induction, note first that $U(\mathfrak{f}_{1})\cong\mathbb{C}[z]$ for every $k$. If $k=1$, then the assertion is given by Lemma 4.3. Assume now that the claim holds for $k-1$. Then the homomorphism in (4.8) can be written as $R\mathop{\#}\mathbb{C}[z]\to S\mathop{\widehat{\#}}\mathfrak{A}_{i_{k}}$, where $R$ and $S$ denote the algebras obtained on the $(k-1)$th step.
Note that $R$ has countable dimension and $g\!:R\to S$ has dense range. Since, in addition, $g$ is a homological epimorphism by the induction hypothesis, Theorem 4.6 implies that $R\mathop{\#}\mathbb{C}[z]\to S\mathop{\widehat{\#}}\mathfrak{A}_{i_{k}}$ is also a homological epimorphism. The claim is proved.
Putting $k=n$ we conclude that $U(\mathfrak{g})\to\widehat{U}(\mathfrak{g})$ is a homological epimorphism. ∎

### v4

@@ 4. Homological epimorphisms and relatively quasi-free algebras
Let $K$ be a Hopf $\mathbin{\widehat{\otimes}}$-algebra, $R$ a $\mathbb{C}[z]$-module algebra of countable dimension and $S$ a $K$$\mathbin{\widehat{\otimes}}$-module algebra. Suppose that $\beta\!:\mathbb{C}[z]\to K$ is a Hopf $\mathbin{\widehat{\otimes}}$-algebra homomorphism and $\alpha\!:R\to S$ is a $\mathbin{\widehat{\otimes}}$-algebra homomorphism that is a $\mathbb{C}[z]$-module morphism. If $\alpha$ and $\beta$ are homological epimorphisms with dense range, then so is
$\alpha\mathop{\widehat{\#}}\beta\!:R\mathop{\#}\mathbb{C}[z]\to S\mathop{\widehat{\#}}K.$
@@
Recall that an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ is said to be relatively quasi-free over $R$ if any admissible singular $R$-extension of $A$ is split; see [26, Definition 7.1] and the discussion therein.
@@
For the proof we need a relative version of Property (UDE). If $A$ is an $R$$\mathop{\widehat{\otimes}}$-algebra and $X$ is an $A$$\mathop{\widehat{\otimes}}$-bimodule, then a derivation $D\!:A\to X$ is called an $R$-derivation if it is an $R$-bimodule morphism (or, equivalently, a left or right $R$-module morphism). The subspace of $\operatorname{Der}(A,X)$ consisting of $R$-derivations is denoted by $\operatorname{Der}_{R}(A,X)$. An $R$-$S$-homomorphism $(f,g)\!:(A,R)\to(B,S)$ induces the linear map
$\operatorname{Der}_{S}(B,X)\to\operatorname{Der}_{R}(A,X)\!:D\mapsto Df$
for every $B$$\mathop{\widehat{\otimes}}$-bimodule $X$; see details in [26].
@@
Since each of $f$ and $g$ has dense range, it suffices to verify (4.3) in the case when $s=g(r)$ and $b=f(a)$ for some $r\in R$ and $a\in A$. Note that
$\eta_{B}(s)b=\eta_{B}(g(r))f(a)=f(\eta_{A}(r))f(a)=f(\eta_{A}(r)a)$
and hence $D^{\prime}(\eta_{B}(s)b)=D(\eta_{A}(r)a)$. Since $D$ is an $R$-derivation, we have $D^{\prime}(\eta_{B}(s)b)=r\cdot D(a)$. On the other hand,
$r\cdot D(a)=f(\eta_{A}(r))\cdot D(a)=\eta_{B}(g(r))\cdot D^{\prime}f(a)=\eta_{B}(s)\cdot D^{\prime}(b).$
@@
The conditions of Proposition 4.7 can be weakened by assuming that $g$ is an epimorphism of $\mathop{\widehat{\otimes}}$-algebras; cf. [26, Proposition 3.8]. But we do not need this variant.
If $A$ is an $R$$\mathop{\widehat{\otimes}}$-algebra, then there is a representing object of the functor $\operatorname{Der}_{R}(A,-)$ from $A\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}A$ to the category of sets. It is denoted by $\Omega^{1}_{R}A$ and called the bimodule of relative differential $1$-forms of $A$; for details see [26, p. 82].
@@
where $\mathbf{h}{{}_{B}}(-,-)$ denotes the vector space of $B$$\mathop{\widehat{\otimes}}$-bimodule morphisms (and similarly for $A$); see [26, eq. (7.3) and p. 83]. Also, consider the natural map
@@
Substituting $\Omega_{S}^{1}B$ for $X$, we obtain a $B$$\mathop{\widehat{\otimes}}$-bimodule morphism
$B\mathop{\widehat{\otimes}}_{A}\Omega^{1}_{R}A\mathop{\widehat{\otimes}}_{A}B\to\Omega_{S}^{1}B$ (4.7)
@@
(cf. [26, Proposition 7.4]) Let $(f,g)\!:(A,R)\to(B,S)$ be an $R$-$S$-homomorphism from an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ to an $S$$\mathop{\widehat{\otimes}}$-algebra $B$. Then the canonical map $\operatorname{Der}_{S}(B,X)\to\operatorname{Der}_{R}(A,X)$ is a bijection for every $B$$\mathop{\widehat{\otimes}}$-bimodule $X$ if and only if the morphism in (4.7) is a $B$$\mathop{\widehat{\otimes}}$-bimodule isomorphism.
@@
Let $X$ be an arbitrary $B$$\mathop{\widehat{\otimes}}$-bimodule. The map in (4.7) is an isomorphism if and only if the map in (4.6) is bijective. It is trivial that the top arrow in (4.4) is bijective if and only if the bottom arrow is bijective and this completes the proof. ∎
@@
By [26, Proposition 7.3], $A$ is relatively quasi-free over $R$ if and only if $\Omega^{1}_{R}A$ is projective in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(A,R)$. Hence the sequence
@@
is a projective resolution in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(A,R)$. Here $\mu_{A}$ is the multiplication on $A$ and $j$ is the bimodule morphism corresponding to the inner derivation
$-\mathop{\mathrm{ad}}\nolimits_{1\otimes 1}\!:A\to A\mathop{\widehat{\otimes}}_{R}A\!:a\mapsto 1\otimes a-a\otimes 1.$
@@
Since $f$ has dense range, it is an epimorphism. Then the canonical map $B\mathop{\widehat{\otimes}}_{A}B\to B$ is a topological isomorphism. Also, since $g$ is an epimorphism, the canonical map $B\mathop{\widehat{\otimes}}_{R}B\to B\mathop{\widehat{\otimes}}_{S}B$ is a topological isomorphism; for both properties of epimorphisms see [26, Proposition 6.1].
We claim that the application the functor $B\mathop{\widehat{\otimes}}_{A}(-)\mathop{\widehat{\otimes}}_{A}B$ to the sequence in (4.8) produces the complex
$0\leftarrow B\leftarrow B\mathop{\widehat{\otimes}}_{S}B\leftarrow\Omega^{1}_{S}B\leftarrow 0$ (4.9)
Indeed, it suffices to show that $B\mathop{\widehat{\otimes}}_{A}\Omega^{1}_{R}A\mathop{\widehat{\otimes}}_{A}B\cong\Omega_{S}^{1}B$. It follows from Proposition 4.7 that the top arrow in (4.4) is bijective for every $X$ and so by Proposition 4.9, we have an isomorphism.
@@
Finally, (4.9) splits in $B\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}S$ by [26, Proposition 7.2]. Hence it also splits in $S\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}S$. This implies that $f$ is a two-sided relative homological epimorphism; see Definition 4.1. ∎
@@
Following [26, Definition 8.3], we say that a projective $(A,R)$$\mathop{\widehat{\otimes}}$-bimodule $P$ satisfies the finiteness condition $(f_{2})$ if it is a retract of a bimodule of the form $A\mathop{\widehat{\otimes}}_{R}R^{n}\mathop{\widehat{\otimes}}_{R}A$ for some $n\in\mathbb{N}$. (We do not need the alternative condition $(f_{1})$ from [26].) Also, an $R$$\mathop{\widehat{\otimes}}$-algebra $A$ is said to be of $(f_{2})$-finite type if in $(A,R)\mbox{-}\!\mathop{\mathsf{mod}}\!\mbox{-}(A,R)$ it has a projective finite-length resolution all of whose bimodules satisfy the finiteness condition $(f_{2})$ [26, Definition 8.4].
Proposition 4.10.
Let $R$ be an associative algebra of countable dimension and $\delta$ a derivation of $R$. Denote the Ore extension $R[z;\delta]$ by $A$. Then the following conditions hold.
(A) $A$ is relatively quasi-free over $R$.
(B) $A$ is an $R$$\mathop{\widehat{\otimes}}$-algebra of $(f_{2})$-finite type.
Proof.
Part (A) is a partial case of Proposition 7.9 in [26]. Moreover, since we deal with an Ore extension with trivial twisting (i.e., the twisting endomorphism is identical), this proposition implies that $\Omega^{1}_{R}A\cong A\mathbin{\otimes_{R}}A$ as an $(A,R)$-bimodule; see also [26, Proposition 7.8]. In particular, $\Omega^{1}_{R}A$ satisfies the finiteness condition $(f_{2})$. Since $A$ is relatively quasi-free by Part (A), it follows from [26, Proposition 7.3.] that $\Omega^{1}_{R}A$ is a projective $(A,R)$-bimodule. Then the sequence in (4.8) is a projective resolution and thus $A$ is of $(f_{2})$-finite type; cf. [26, Example 8.1]. This completes the proof of Part (B). ∎
The following result gives a direct relation between homological epimorphism and Property (UDE).
Theorem 4.11.
Every homological epimorphism of $\mathop{\widehat{\otimes}}$-algebras satisfies Property (UDE).
Proof.
Let $\varphi\!:A\to B$ be a homological epimorphism. Consider the (unnormalized) bimodule bar resolution of $A$ [17, Section III.2.3]. It has the form
$0\leftarrow A\xleftarrow{\mu_{A}}A\mathop{\widehat{\otimes}}A\xleftarrow{d_{1}}A\mathop{\widehat{\otimes}}A\mathop{\widehat{\otimes}}A\xleftarrow{d_{2}}\cdots\leftarrow A^{\mathop{\widehat{\otimes}}n}\leftarrow\cdots,$ (4.10)
where $\mu_{A}$ is the multiplication map and $d_{n}$, $n\in\mathbb{N}$ are is given by standard formulas. Applying $B\mathop{\widehat{\otimes}}_{A}(-)\mathop{\widehat{\otimes}}_{A}B$ to this resolution we get the sequence
$0\leftarrow B\xleftarrow{\mu_{B}}B\mathop{\widehat{\otimes}}B\xleftarrow{\widetilde{d}_{1}}B\mathop{\widehat{\otimes}}A\mathop{\widehat{\otimes}}B\xleftarrow{\widetilde{d}_{2}}\cdots,$
which is admissible since $\varphi$ is a homological epimorphism. Hence $\operatorname{Coker}\widetilde{d}_{2}\to\operatorname{Ker}\mu_{B}$ is a topological isomorphism.
Note that $\operatorname{Ker}\mu_{A}\cong\Omega^{1}A$ and $\operatorname{Ker}\mu_{B}\cong\Omega^{1}B$ [26, Proposition 7.2] and, moreover, $\operatorname{Coker}d_{2}\cong\Omega^{1}A$ because (4.10) is admissible. Since the functor of relative tensor product preserves cokernels [27, Proposition 3.3], we have that $B\mathop{\widehat{\otimes}}_{A}\Omega^{1}A\mathop{\widehat{\otimes}}_{A}B\to\Omega^{1}B$ defined in (4.7) is a $B$$\mathop{\widehat{\otimes}}$-bimodule isomorphism. Applying Proposition 4.9 with $R=S=\mathbb{C}$, we conclude that $\varphi$ satisfies Property (UDE). ∎
@@
Recall that we denote by $\beta$ the Hopf $\mathop{\widehat{\otimes}}$-algebra homomorphism $\mathbb{C}[z]\to K$. Put $A=R\mathop{\#}\mathbb{C}[z]$, $B=S\mathop{\widehat{\#}}K$, $g=\alpha$ and $f=\alpha\mathop{\widehat{\#}}\beta$ (the latter is well defined by Lemma 2.3 since $\alpha$ is a $\mathbb{C}[z]$-module morphism). It suffices to check the conditions in Theorem 1.1, i.e.,
@@
Condition (2) follows from Theorem 4.6. Indeed, since $\alpha$ and $\beta$ are homological epimorphisms, they satisfy Property (UDE) by Theorem 4.11. Furthermore, Theorem 3.5 implies that $f$ also satisfies Property (UDE). Since $R$ is a $\mathbb{C}[z]$-module algebra, we have the action is induced by a derivation. Therefore $R\mathop{\#}\mathbb{C}[z]$ is an Ore extension (see Remark 2.2) and hence it is relatively quasi-free by Part (A) of Proposition 4.10. Thus all the conditions in Theorem 4.6 hold and therefore $f$ is a two-sided relative homological epimorphism. Finally, note that every two-sided relative homological epimorphism is a one-sided relative homological epimorphism [26, Proposition 6.5].
To verify Condition (3) note that $S\mathop{\widehat{\#}}K$ is isomorphic to $S\mathop{\widehat{\otimes}}K$ as an $S$$\mathop{\widehat{\otimes}}$-module and so it is projective being a free module (see, e.g., [17, Chapter III, Proposition 1.25]). Similarly, $R\mathop{\#}\mathbb{C}[z]$ is a projective $R$-module.
Finally, Condition (4) holds by Part (B) of Proposition 4.10. ∎
@@
For the necessity see [24, Theorem 3.6].
To show the sufficiency consider the decomposition in (2.9) from Theorem 2.5. We claim that the truncated homomorphism
@@
Proceeding by induction, note first that $U(\mathfrak{f}_{k})\cong\mathbb{C}[z]$ for every $k$. If $k=1$, then the assertion is given by Proposition 4.4.
Assume now that the claim holds for $k-1$. Then the homomorphism in (4.11) can be written as
$\alpha\mathop{\widehat{\#}}\beta\!:R\mathop{\#}\mathbb{C}[z]\to S\mathop{\widehat{\#}}\mathfrak{A}_{i_{k}},$
where $R$ and $S$ denote the algebras obtained on the $(k-1)$th step. Here $S$ is a $\mathfrak{A}_{i_{k}}$$\mathbin{\widehat{\otimes}}$-module algebra, $\beta\!:\mathbb{C}[z]\to\mathfrak{A}_{i_{k}}$ is a Hopf $\mathbin{\widehat{\otimes}}$-algebra homomorphism and $\alpha\!:R\to S$ is a $\mathbin{\widehat{\otimes}}$-algebra homomorphism that is a $\mathbb{C}[z]$-module morphism.
Note that $R$ has countable dimension and each of $\alpha$ and $\beta$ has dense range. Moreover, $\alpha$ and $\beta$ are homological epimorphisms (the first by the induction hypothesis and the second by Proposition 4.4). Thus all the assumptions in Theorem 4.5 hold and hence $R\mathop{\#}\mathbb{C}[z]\to S\mathop{\widehat{\#}}\mathfrak{A}_{i_{k}}$ is also a homological epimorphism. The claim is proved.
Finally, putting $k=n$ we conclude that $U(\mathfrak{g})\to\widehat{U}(\mathfrak{g})$ is a homological epimorphism. ∎